OpenStudy (anonymous):

Im working on factoring trinomials. Can anyone tell me how the factors of a quadratic equation can help you find the x intercepts?

OpenStudy (anonymous):

x intercepts means where it crosses the x - axis. on the x axis the y value is 0. so it is the same as saying 'set your equation = 0 and solve' if i have \[y=x^2-x-2\] and i want to know where it crosses the x - axis that is the same as saying solve for x: \[x^2-x-2=0\] now if i can factor this is easy: \[x^2-x-2=0\] \[(x-2)(x+1)=0\] \[x=2\]or \[x=-1\] and therefore the x intercepts are 2 and -1 that is why i needed to factor, to find out where it is 0.